Following a heart attack, a physician orders a patient to begin taking a beta-blocker to treat angina and hypertension. The tablets are available in 50 mg. If the physician gradually increases the patient’s dosage, calculate the number of tablets required if the doctor requests 150 mg/day, 200 mg/day and 225 mg/day


1
Expert's answer
2021-10-28T12:15:02-0400

If the doctor requests 150 mg/day: "3 \times 50 = 150 \;mg"

3 tablets per day

If the doctor requests 200 mg/day: "4 \times 50 = 200 \;mg"

4 tablets per day

If the doctor requests 225 mg/day: "4.5 \times 50 = 225 \;mg"

4.5 tablets per day